# Dynamic linear regression models for forecasting time series with semi long memory errors

### v1

@@ Abstract.
Dynamic linear regression models forecast the values of a time series based on a linear combination of a set of exogenous time series while incorporating a time series process for the error term. This error process is often assumed to follow an autoregressive integrated moving average (ARIMA) model, or seasonal variants thereof, which are unable to capture a long-range dependency structure of the error process. We propose a novel dynamic linear regression model that incorporates the long-range dependency feature of the errors. We demonstrate that the proposed error process may (i) have a significant impact on the posterior uncertainty of the estimated regression parameters and (ii) improve the model’s forecasting ability. We develop a Markov chain Monte Carlo method to fit general dynamic linear regression models based on a frequency domain approach that enables fast, asymptotically exact Bayesian inference for large datasets. We demonstrate that our approximate algorithm is faster than the traditional time domain approaches, such as the Kalman filter and the multivariate Gaussian likelihood, while retaining a high accuracy when approximating the posterior distribution. We illustrate the method in simulated examples and two energy forecasting applications.
Keywords: Dynamic linear regression, forecasting, Bayesian inference.
@@ 1. Introduction
Forecasting time series data plays an important role in various fields, such as engineering, economics, and climate sciences. Forecasting a single output time series may be more accurate by using a linear combination of one or more exogenous time series that explains some of its historical variation (Hyndman and Athanasopoulos,, 2018). However, the linear combination of the exogenous time series often does not capture all the serial correlation present in the output time series, resulting in errors that are autocorrelated, i.e. serially correlated. Dynamic linear regression models (Pankratz,, 2012) provide a framework that relates the output time series to a linear combination of the exogenous time series and models the resulting error term as a time series process to account for serially correlated errors. Dynamic linear regression (DLR) models are a particular case of the more general class transfer function models (Box et al.,, 2015).
@@
The standard dynamic linear regression model assumes that the error process, i.e. the time series process for the error terms, follows an autoregressive integrated moving average process (ARIMA). This approach allows the error process to have an autoregressive component and a moving average component for the white noise error. While this error process accommodates a wide range of stationary processes, it may require a large number of autoregressive components to accommodate error processes that show significant autocorrelation between distant time periods. To model the autocovariance of the error process parsimoniously, we use the autoregressive tempered fractionally integrated moving average (ARTFIMA) (Meerschaert et al.,, 2014; Sabzikar et al.,, 2019). The ARTFIMA class nests the well-known autoregressive fractionally integrated moving average (ARFIMA) (Granger and Joyeux,, 1980) model, which is useful for modelling time series with so-called long memory.
@@
The ARTFIMA model has the following three advantages compared to the ARFIMA model, where the latter two are of particular importance for the frequency domain estimation approach for dynamic linear regression models proposed in this paper. First, the autocovariance function of the ARFIMA model decays at such a slow rate (power-law decay) that it is not absolutely summable, making it hard to analyse (Sabzikar et al.,, 2019). Such a process is referred to as having long memory. In contrast, the ARTFIMA model has a summable autocovariance function, which exhibits long-range dependence for a number of lags but eventually decays exponentially fast (i.e. much faster than power law). Such a process is referred to as having semi-long memory. Second, the spectral density of the ARFIMA process diverges as the frequency approaches zero. In practice, however, the empirical power spectrum (the estimate of the spectral density) is bounded for small frequencies. This is empirically illustrated in Meerschaert et al., (2014); Sabzikar et al., (2019), who show that the ARTFIMA process (whose spectral density does not diverge as the frequency approaches zero) fits the power spectrum better than the ARFIMA process for low frequencies; see also Figure 10 showing this carries over to the DLR case. This stylised fact is important for our purpose as we develop an estimation method based on a parametric Whittle likelihood (Whittle,, 1953), which is a frequency domain approximation of the time domain likelihood that represents the information in the data via the spectral density. Third, the Whittle approximation fails to hold for long memory processes, especially for small frequencies (Robinson,, 1995; Rousseau et al.,, 2012). Section 4.2 demonstrates this drawback of ARFIMA models via a simulation study and also shows that the ARTFIMA model overcomes this issue. We show that resorting to the time domain likelihood for estimating a dynamic linear regression with ARFIMA errors (Doornik and Ooms,, 2003) is computationally much more costly than our approach of a frequency domain likelihood with ARTFIMA errors. Moreover, in terms of prediction accuracy, our model is often superior, or at least on par with.
@@
The model parameters in dynamic linear regression models consist of the regression coefficients that form the linear combination of the exogenous time series and the parameters of the error process. For the standard dynamic linear regression models with ARIMA errors that are normally distributed, efficient likelihood-based inference can be carried out by finding the finite-dimensional state space representation of the model and using the Kalman filter to integrate out the unobserved time-varying error terms. This may still be computationally costly with many time observations, especially in Bayesian inference, which typically requires many posterior samples for reliable inference. For each such sample, the Kalman filter needs to cycle through all observations, which can be prohibitively expensive, especially for large time series. The Whittle log-likelihood can be derived using large sample properties of the so-called periodogram data, which are formed via the discrete Fourier transform of the time domain data. The Whittle log-likelihood is directly a function of the regression coefficients and the error process parameters without needing to integrate out unobserved error terms. However, in dynamic linear regression models, the periodogram data becomes a function of the regression coefficients, which requires recomputing the discrete Fourier transform in every iteration. We show how this can be circumvented, and thus, our algorithm requires computing the discrete Fourier transform only once before running the sampling algorithm, thereby obtaining significant computational gains compared to the time domain log-likelihood based on the Kalman filter. Our frequency domain approach is also applicable to a dynamic linear regression model with ARTFIMA errors, where a finite-dimensional state space representation is not readily available.
@@
To summarise, our article has two important contributions. First, we propose a frequency domain estimation approach for dynamic linear models that significantly outperforms estimation approaches based on the time domain likelihood in terms of computing time, especially when a finite-dimensional state space representation of the model is not available. Second, we utilise a semi-long memory process for the error process and show that it may provide more accurate forecasts compared to both the standard dynamic linear model and that of using a long memory process for the error term, especially for datasets with a complex long-range dependency. Moreover, we demonstrate that the semi-long memory process may significantly impact the posterior distribution of the regression parameter.
@@
The rest of the article is organised as follows. Section 2 reviews existing dynamic linear models and presents our extension. Section 3 introduces the necessary frequency domain tools, outlines our estimation method, and validates its performance relative to the time domain likelihood. Section 5 presents applications for two real-world electricity demand datasets.
@@ 2.1. Standard dynamic linear regression models
Let $\boldsymbol{X}_{t}=(X_{1t},\dots,X_{mt})^{\top}\in\mathbb{R}^{m}$ be a set of $m$ exogenous stationary time series observed at time $t$. A dynamic linear regression model models the single output time series $Y_{t}\in\mathbb{R}$ as a linear combination of the exogenous $\boldsymbol{X}_{t}$, i.e.
(2.1) $Y_{t}$ $=\boldsymbol{X}_{t}^{\top}\boldsymbol{\beta}+\eta_{t},$
@@
When the error process $\eta_{t}$ does not show any temporal dependence, (2.1) is a standard linear regression (given it has a constant variance) and can easily be estimated using standard linear regression approaches (or modified versions thereof if heteroscedasticity is present). However, in many applications, the unobserved factors are time-varying, resulting in serially correlated errors $\eta_{t}$. The standard dynamic linear regression model assumes that $\eta_{t}$ is an autoregressive integrated moving average process, denoted ARIMA$(p,d,q)$ and defined as
(2.2) $\phi_{p}(B)\Delta^{d}\eta_{t}$ $=\psi_{q}(B)\varepsilon_{t},$
@@
(2.3) $\phi_{p}(B)\phi_{P}^{\star}(B^{s})\Delta^{d}\Delta^{D}_{s}\eta_{t}=\psi_{q}(B)\psi_{Q}^{\star}(B^{s})\varepsilon_{t},$
@@
To carry out inference in (2.1), one typically assumes that $\eta_{t}$ follows a normal distribution. The resulting log-likelihood is then a multivariate Gaussian distribution with a Toeplitz covariance matrix given that $\eta_{t}$ is stationary (Doornik and Ooms,, 2003). Inversion of such a matrix is usually performed via the Levinson-Durbin algorithm (Levinson,, 1946; Durbin,, 1960) in $\mathcal{O}(T^{2})$ operations. However, recent approaches, known as superfast Toeplitz algorithms, can solve the matrix system in $\mathcal{O}(T\log_{2}^{2}(T))$ operations, which becomes more efficient than the Levinson-Durbin algorithm when $T>256$ (Ammar and Gragg,, 1988). A method that scales even better is to find the finite-dimensional state space representation of the model in (2.1) with the error process following either (2.2) or (2.3), in which the resulting likelihood can be evaluated in $\mathcal{O}(T)$ using the Kalman filter.
@@ 2.2. Long memory dynamic linear regression models
In many applications, a pair of observations separated by a long time interval exhibit a non-negligible correlation. The standard dynamic linear regression model with the error process in (2.2) has an autocovariance function whose absolute value decays exponentially fast and can thus not capture this feature. A common approach to model time series with long memory is to consider so-called fractional differencing.
@@
(2.4) $\phi_{p}(B)\Delta^{d}\eta_{t}$ $=\psi_{q}(B)\varepsilon_{t},$
where $\phi_{p}(B)$ and $\psi_{q}(B)$ are defined as in (2.2). For $d\notin\mathbb{Z}$ the fractional differencing operator is defined via the fractional binomial theorem
@@
Doornik and Ooms, (2003) show that maximum likelihood estimation for ARFIMA models can be carried out in $\mathcal{O}(T^{2})$ time, and this also applies for the model in (2.1) with the error process following (2.4). Chan and Palma, (1998) show that there is no finite-dimensional state space representation of an ARFIMA process. They propose an infinite-dimensional state space representation for which the Kalman filter can be computed in a finite number of steps equal to $T$; however, the resulting computation is $\mathcal{O}(T^{3})$.
2.3. Semi long memory dynamic linear regression models
The autoregressive tempered fractional integrated moving-average (ARTFIMA) (Meerschaert et al.,, 2014; Sabzikar et al.,, 2019) is an extension of the ARFIMA model that incorporates a tempering parameter $\lambda$. The model for the error process is then
(2.5) $\phi_{p}(B)\Delta^{d,\lambda}\eta_{t}$ $=\psi_{q}(B)\varepsilon_{t},$
@@ 2.3. Semi long memory dynamic linear regression models
(2.6) $=\sum_{j=0}^{\infty}(-1)^{j}\frac{\Gamma(1+d)}{\Gamma(1+d-j)j!}\exp(-\lambda j)\eta_{t-j}.$
Note that when $\lambda=0$, the ARTFIMA model becomes a stationary ARFIMA model if $-0.5<d<0.5$, provided the roots of the polynomial $\phi_{p}(z)$ are outside the unit circle in the complex plane. When $\lambda=0$ and $d$ is integer-valued, the ARTFIMA model becomes an ARIMA model. The ARTFIMA model is stationary for all $d\notin\mathbb{Z}$ and $\lambda>0$ provided that the root condition of the polynomial $\phi_{p}(z)$ is satisfied (Sabzikar et al.,, 2019). The ARTFIMA process has several advantages over the ARFIMA process, which are outlined in Section 1.
Similar to the ARFIMA model, a finite-dimensional state space representation that allows efficient Kalman filter computations is not readily available. Instead, the log-likelihood can be computed using a multivariate Gaussian distribution with a Toeplitz covariance matrix formed via the autocovariance function in Sabzikar et al., (2019, Theorem 2.5 (b)). However, a more computationally efficient approach is forming a frequency domain log-likelihood for the so-called periodogram data described in Section 3. The approach relies on the spectral density, also known as the power, of the process which we now describe.
Let $\omega$ denote the angular frequency. The spectral density $f(\omega)$ is the Fourier transform of the covariance function and shows how the variance of the time series is distributed among the frequencies $-\pi\leq\omega\leq\pi$. The spectral density when $\eta_{t}$ follows the ARTFIMA process in (2.5) is (Sabzikar et al.,, 2019, Theorem 2.5 (a))
(2.7) $f(\omega;\boldsymbol{\vartheta})$ $=\frac{\sigma_{\varepsilon}^{2}}{2\pi}\left|1-e^{-(\lambda+i\omega)}\right|^{-2d}\left|\frac{\psi_{q}(e^{-i\omega})}{\phi_{p}(e^{-i\omega})}\right|^{2},$
@@
(2.8) $\phi_{p}(B)\phi_{P}^{\star}(B^{s})\Delta^{d,\lambda}\Delta^{D}_{s}\eta_{t}=\psi_{q}(B)\psi_{Q}^{\star}(B^{s})\varepsilon_{t},$
where all quantities have been previously defined, has the spectral density
(2.9) $f(\omega;\boldsymbol{\vartheta})$ $=\frac{\sigma_{\varepsilon}^{2}}{2\pi}\left|1-e^{-(\lambda+i\omega)}\right|^{-2d}\left|\frac{\psi_{q}(e^{-i\omega})\psi^{\star}_{Q}(e^{-is\omega})}{\phi_{p}(e^{-i\omega})\phi^{\star}_{P}(e^{-is\omega})}\right|^{2}.$
Figure 1 shows the spectral density for both the ARFIMA and ARTFIMA process for different values of the fractional differencing parameter $d$. An important difference that we have stressed is the limiting behaviour as $\omega\rightarrow 0$, where the spectral density of the ARFIMA process diverges when $d>0$, whereas that of the ARTFIMA process does not. When estimating the power for real data, it is often observed to be bounded. Hence the ARTFIMA model provides a better fit for small frequencies; see for example Figure 10.
Figure 1. Spectral densities of different ARFIMA and ARTFIMA models with $p=1$ and $q=0$. For both processes, the parameter $\phi=0.5$ and ARTFIMA $\lambda=0.045$.

### v2

# Dynamic linear regression models for forecasting time series with semi-long memory errors

@@ Abstract.
Dynamic linear regression models forecast the values of a time series based on a linear combination of a set of exogenous time series while incorporating a time series process for the error term. This error process is often assumed to follow a stationary autoregressive integrated moving average (ARIMA) model, or its seasonal variants, which is unable to capture a long-range dependence structure (long memory) of the error process. We propose a novel dynamic linear regression model that incorporates the long-range dependence feature of the errors and show that the proposed error process may: (i) have a significant impact on the posterior uncertainty of the estimated regression parameters and (ii) improve the model’s forecasting ability. We develop a Markov chain Monte Carlo method to fit general dynamic linear regression models based on a frequency domain approach that enables fast, asymptotically exact Bayesian inference for large datasets. We demonstrate that our approximate algorithm is faster than the traditional time domain approaches, such as the Kalman filter and the multivariate Gaussian likelihood, while producing a highly accurate approximation to the posterior. The method is illustrated in simulated examples and two energy forecasting applications.
Keywords: Bayesian inference, Frequency domain methods, Energy forecasting.
@@ 1. Introduction
Forecasting time series data plays an important role in various fields, such as engineering, economics, and climate sciences. Forecasting a single output time series may be more accurate by using linear combinations of exogenous variables that explain some of its historical variation (Hyndman and Athanasopoulos,, 2021). However, the linear combination of the exogenous time series often does not capture all the serial correlation present in the output time series, resulting in autocorrelated errors. Dynamic linear regression models (Pankratz,, 2012) provide a framework that relates the output time series to a linear combination of the exogenous time series and models the resulting error term as a time series process to account for serially correlated errors. Dynamic linear regression (DLR) models are a particular case of the more general class of transfer function models (Box et al.,, 2015).
@@
The standard dynamic linear regression model assumes that the error process follows an autoregressive integrated moving average process (ARIMA). This approach allows the error process to have an autoregressive component and a moving average component for its white noise error term. While this error process accommodates a wide range of stationary processes, it requires a large number of autoregressive components to accommodate error processes that show significant autocorrelation between distant time periods. To capture such autocorrelations of the error process parsimoniously, we employ the autoregressive tempered fractionally integrated moving average (ARTFIMA) (Meerschaert et al.,, 2014; Sabzikar et al.,, 2019). The ARTFIMA class nests the well-known autoregressive fractionally integrated moving average (ARFIMA) (Granger and Joyeux,, 1980) model, which is useful for time series with so-called long memory.
@@
The ARTFIMA model has the following three advantages compared to the ARFIMA model, where the latter two are particularly important for the frequency domain estimation approach for dynamic linear regression models proposed in this paper. First, the autocovariance function of the ARFIMA model decays slowly as a power law (Granger and Joyeux,, 1980) which is not absolutely summable. Such a process is referred to as having long memory (McLeod and Hipel,, 1978). In contrast, the ARTFIMA model is appealing since its autocovariance function exhibits long-range dependence for several lags, but eventually decays exponentially fast (i.e. much faster than power law) and is hence absolutely summable. Such a process is referred to as having semi-long memory and is easier to analyse (Sabzikar et al.,, 2019). Second, the spectral density, also known as the power, of the ARFIMA process diverges as the frequency approaches zero. In practice, however, a stylised fact of the empirical power spectrum (the estimate of the spectral density) is that it is bounded for low frequencies. This is illustrated in Meerschaert et al., (2014) and Sabzikar et al., (2019), who show that the ARTFIMA process (whose spectral density does not diverge as the frequency approaches zero) fits the power spectrum better than the ARFIMA process for low frequencies; see also Figure 9 showing this carries over to the DLR case. This stylised fact is important for our purpose as we develop an estimation method based on a parametric Whittle likelihood (Whittle,, 1953), which is a frequency domain approximation of the time domain likelihood that represents the information in the data via the spectral density. Third, the Whittle approximation fails to hold for long memory processes, especially for small frequencies (Robinson,, 1995; Rousseau et al.,, 2012). Section 4.2 demonstrates this drawback for DLR models based on ARFIMA errors via a simulation study and also shows that ARTFIMA errors overcomes this issue. We show that resorting to the time domain likelihood for estimating a dynamic linear regression with ARFIMA errors (Doornik and Ooms,, 2003) is computationally much more costly than our frequency domain approach. Moreover, in terms of prediction accuracy, our model is often superior, or at least on par with DLR with ARFIMA and ARMA errors.
@@
The parameters in dynamic linear regression models consist of the regression coefficients that form the linear combination of the exogenous time series and the parameters of the error process. For the standard dynamic linear regression models with ARIMA errors that are normally distributed, efficient likelihood-based inference can be carried out by finding the finite-dimensional state space representation of the model and using the Kalman filter to integrate out the unobserved time-varying error terms (Anderson and Moore,, 2005). This may still be computationally costly with many time observations, especially in Bayesian inference, which typically requires many posterior samples for reliable inference. For each such sample, the Kalman filter needs to cycle through all the observations, which can be prohibitively expensive, especially for large time series. The Whittle log-likelihood can be derived using large sample properties of the so-called periodogram data, which are formed via the discrete Fourier transform of the time domain data. The Whittle log-likelihood is directly a function of the regression coefficients and the error process parameters without needing to integrate out unobserved error terms. However, in DLR models, the periodogram data becomes a function of the regression coefficients, which requires recomputing the discrete Fourier transform in every iteration. We show how this can be avoided, and thus our algorithm requires computing the discrete Fourier transform only once before running the sampling algorithm, thereby obtaining significant computational gains compared to the time domain log-likelihood based on the Kalman filter. Our frequency domain approach also applies to a dynamic linear regression model with ARTFIMA errors, where a finite-dimensional state space representation is not readily available.
@@
To summarise, our article makes two important contributions. First, we propose a frequency domain estimation approach for dynamic linear models that significantly outperforms estimation approaches based on the time domain likelihood in terms of computing time, especially when a finite-dimensional state space representation of the model is unavailable. Second, we utilise a semi-long memory process for the error process and show that it may provide more accurate forecasts compared to both the standard dynamic linear model and that of using a long memory process for the error term, especially for datasets with complex long-range dependence. Moreover, we demonstrate that the semi-long memory process may significantly impact the posterior distribution of the regression parameters.
@@
The rest of the article is organised as follows. Section 2 reviews existing dynamic linear models and presents our extension. Section 3 introduces the necessary frequency domain tools, outlines our estimation method, and validates its performance relative to the time domain likelihood. Section 4 includes relevant simulation studies. Section 5 presents applications for two real-world electricity demand datasets.
@@ 2.1. Standard dynamic linear regression models
Let $\boldsymbol{X}_{t}=(X_{1t},\dots,X_{mt})^{\top}\in\mathbb{R}^{m}$ be a set of $m$ exogenous stationary time series observed at time $t$. A dynamic linear regression models the scalar output time series $Y_{t}\in\mathbb{R}$ as a linear combination of the exogenous $\boldsymbol{X}_{t}$, i.e.
$Y_{t}$ $=\boldsymbol{X}_{t}^{\top}\boldsymbol{\beta}+\eta_{t},$ (2.1)
@@
When the error process $\eta_{t}$ is an independent series, (2.1) is a standard linear regression (if it has a constant variance) and can easily be estimated using standard linear regression approaches (or modified versions thereof if heteroscedasticity is present). However, in many applications, the unobserved factors are time-varying, resulting in serially correlated errors $\eta_{t}$. The standard dynamic linear regression model assumes that $\eta_{t}$ is an autoregressive integrated moving average process, denoted ARIMA$(p,d,q)$ and defined as
$\phi_{p}(B)\Delta^{d}\eta_{t}$ $=\psi_{q}(B)\varepsilon_{t},$ (2.2)
@@
$\phi_{p}(B)\phi_{P}^{\star}(B^{s})\Delta^{d}\Delta^{D}_{s}\eta_{t}=\psi_{q}(B)\psi_{Q}^{\star}(B^{s})\varepsilon_{t},$ (2.3)
@@
To carry out inference in (2.1), it is typically assumed that $\eta_{t}$ is normally distributed. The resulting log-likelihood is then a multivariate Gaussian distribution with a Toeplitz covariance matrix given that $\eta_{t}$, or after integer differencing of $\eta_{t}$, is stationary (Doornik and Ooms,, 2003). Inversion of such a matrix is usually performed via the Levinson-Durbin algorithm (Levinson,, 1946; Durbin,, 1960) in $\mathcal{O}(T^{2})$ operations. However, recent approaches, known as superfast Toeplitz algorithms, can solve the matrix system in $\mathcal{O}(T\log_{2}^{2}(T))$ operations, which becomes more efficient than the Levinson-Durbin algorithm when $T>256$ (Ammar and Gragg,, 1988). A method that scales even better is to find the finite-dimensional state space representation of the model in (2.1) with the error process following either (2.2) or (2.3), in which the resulting likelihood can be evaluated in $\mathcal{O}(T)$ using the Kalman filter.
@@ 2.2. Long memory dynamic linear regression models
In many applications, a pair of observations separated by a long time interval exhibit a non-negligible correlation. The standard dynamic linear regression model with the error process in(2.2) has an autocovariance function whose absolute value decays exponentially fast and thus cannot capture this feature. A common approach to model time series with long memory is to consider so-called fractional differencing.
@@
$\phi_{p}(B)\Delta^{d}\eta_{t}$ $=\psi_{q}(B)\varepsilon_{t},$ (2.4)
where $\phi_{p}(B)$ and $\psi_{q}(B)$ are defined as in (2.2). The fractional differencing operator is defined via the fractional binomial theorem
@@
Doornik and Ooms, (2003) show that maximum likelihood estimation for ARFIMA models is possible in $\mathcal{O}(T^{2})$ time, and this also applies for the model in (2.1) with the error process following (2.4). Chan and Palma, (1998) show that there is no finite-dimensional state space representation of an ARFIMA process. They propose an infinite-dimensional state space representation for which the Kalman filter can be computed in $T$ steps; however, the resulting computation is $\mathcal{O}(T^{3})$.
2.3. Semi-long memory dynamic linear regression models
The autoregressive tempered fractional integrated moving-average (Meerschaert et al.,, 2014; Sabzikar et al.,, 2019) (ARTFIMA) extends the ARFIMA model, incorporating a tempering parameter $\lambda$. The model for the error process is then
$\phi_{p}(B)\Delta^{d,\lambda}\eta_{t}$ $=\psi_{q}(B)\varepsilon_{t},$ (2.5)
@@ 2.3. Semi-long memory dynamic linear regression models
$=\sum_{j=0}^{\infty}(-1)^{j}\frac{\Gamma(1+d)}{\Gamma(1+d-j)j!}\exp(-\lambda j)\eta_{t-j}.$ (2.6)
Note that when $\lambda=0$, the ARTFIMA model is a stationary ARFIMA model if $-0.5<d<0.5$, provided the roots of the polynomial $\phi_{p}(z)$ are outside the unit circle. When $\lambda=0$ and $d$ is an integer, the ARTFIMA model becomes an ARIMA model. The ARTFIMA model is stationary for all $d\notin\mathbb{Z}$ and $\lambda>0$ provided that the root condition of the polynomial $\phi_{p}(z)$ is satisfied (Sabzikar et al.,, 2019). Section 1 outlines the advantages the ARTFIMA process has over ARFIMA.
Similarly to the ARFIMA model, a finite-dimensional state space representation that allows efficient Kalman filter computations is unavailable for the ARTFIMA model. Instead, the log-likelihood can be computed using a multivariate Gaussian distribution with a Toeplitz covariance matrix formed via the autocovariance function in Sabzikar et al., (2019, Theorem 2.5 (b)). However, Section 3 describes a more computationally efficient approach based on a frequency domain log-likelihood for the periodogram data. The approach relies on the spectral density of the process, which we now describe.
Let $\omega$ denote an angular frequency. The spectral density $f(\omega)$ is the Fourier transform of the covariance function and shows how the variance of the time series is distributed among the frequencies $-\pi\leq\omega\leq\pi$. The spectral density when $\eta_{t}$ follows the ARTFIMA process in (2.5) is (Sabzikar et al.,, 2019, Theorem 2.5 (a))
$f(\omega;\boldsymbol{\vartheta})$ $=\frac{\sigma_{\varepsilon}^{2}}{2\pi}\left|1-e^{-(\lambda+\mathrm{i}\omega)}\right|^{-2d}\left|\frac{\psi_{q}(e^{-\mathrm{i}\omega})}{\phi_{p}(e^{-\mathrm{i}\omega})}\right|^{2},$ (2.7)
@@
$\phi_{p}(B)\phi_{P}^{\star}(B^{s})\Delta^{d,\lambda}\Delta^{D}_{s}\eta_{t}=\psi_{q}(B)\psi_{Q}^{\star}(B^{s})\varepsilon_{t},$ (2.8)
where all quantities have been previously defined. The spectral density of $\Delta^{D}_{s}\eta_{t}$ in (2.8) is
$f(\omega;\boldsymbol{\vartheta})$ $=\frac{\sigma_{\varepsilon}^{2}}{2\pi}\left|1-e^{-(\lambda+\mathrm{i}\omega)}\right|^{-2d}\left|\frac{\psi_{q}(e^{-\mathrm{i}\omega})\psi^{\star}_{Q}(e^{-\mathrm{i}s\omega})}{\phi_{p}(e^{-\mathrm{i}\omega})\phi^{\star}_{P}(e^{-\mathrm{i}s\omega})}\right|^{2}.$ (2.9)
Figure 1 shows the spectral density for both the ARFIMA and ARTFIMA process for different values of the fractional differencing parameter $d$. An important difference that we have stressed is the limiting behaviour as $\omega\rightarrow 0$, where the spectral density of the ARFIMA process diverges when $d>0$, whereas that of the ARTFIMA process does not. When estimating the power for real data, it is often observed to be bounded as discussed in Section 1. Hence, the ARTFIMA model provides a better fit for small frequencies; see, for example, Figure 9.
Figure 1. Spectral densities of ARFIMA and ARTFIMA models with $p=1$ and $q=0$. For both processes, $\phi=0.5$ and ARTFIMA $\lambda=0.045$.